# Time-Dependent VAE for Building Latent Factor from Visual Neural Activity with Complex Dynamics

### v1

@@ Abstract
Seeking high-quality neural latent representations to reveal the intrinsic correlation between neural activity and behavior or sensory stimulation has attracted much interest. Currently, some deep latent variable models rely on behavioral information (e.g., movement direction and position) as an aid to build expressive embeddings while being restricted by fixed time scales. Visual neural activity from passive viewing lacks clearly correlated behavior or task information, and high-dimensional visual stimulation leads to intricate neural dynamics. To cope with such conditions, we propose Time-Dependent SwapVAE, following the approach of separating content and style spaces in Swap-VAE, on the basis of which we introduce state variables to construct conditional distributions with temporal dependence for the above two spaces. Our model progressively generates latent variables along neural activity sequences, and we apply self-supervised contrastive learning to shape its latent space. In this way, it can effectively analyze complex neural dynamics from sequences of arbitrary length, even without task or behavioral data as auxiliary inputs. We compare TiDe-SwapVAE with alternative models on synthetic data and neural data from mouse visual cortex. The results show that our model not only accurately decodes complex visual stimuli but also extracts explicit temporal neural dynamics, demonstrating that it builds latent representations more relevant to visual stimulation.
@@ 1 Introduction
With the rapid development of neural recording technologies, researchers are now able to simultaneously record the spiking activity of large populations of neurons, providing new avenues for exploring the brain [29]. This type of data has sparked strong interest in proposing new methods to uncover the underlying dynamical structure of neural population activity related to brain functions, usually called latent variable methods [31, 30, 23]. These methods aim to capture the variability in simultaneously recorded neural population data with a limited number of latent variables, providing insights into neural representation, and its relationship with cognition or behavior [10, 15, 18]. Popular methods include principal component analysis [28, 3, 6], factor analysis [34, 24, 11, 9], Gaussian process [32, 35], and variational models [13, 22, 36, 19]. Most of these studies have dealt with neural population data recorded from the motor system [34, 6, 22, 36, 19, 25], while only a few have dealt with data from the visual system [11, 25], or other brain areas [28, 3].
Passive viewing of visual stimuli is an important and commonly adopted experimental paradigm, especially for exploring visual coding and processing [27]. Most latent variable models for studying visual neural activity are limited to simple stimuli, e.g., drifting gratings, and encounter some constraints for complex natural stimuli, such as imposing a strong prior assumption about the dynamics of the latent space [11, 13, 35] and processing each time window independently [21]. Recently, to deal with neural activity with complex dynamics under movie stimuli, Schneider et al. proposed CEBRA, producing consistent and high-performance latent representations using fixed temporal filters [25]. Yet, effectively exploiting the temporal dependence of variable-length sequences to discover latent manifolds, in the absence of task or behavioral information, is still challenging.
@@
To address this problem, we innovatively propose Time-Dependent SwapVAE (TiDe-SwapVAE), which utilizes state variables to construct conditional distributions for content and style latent variables with temporal dependence. We evaluate our model on synthetic and mouse visual datasets. We show that our model builds meaningful latent representations that are highly correlated with complex visual stimuli, which gives new insights into the intrinsic relationship between neural activity and visual stimulation. Specifically, our main contributions are as follows.
@@
  • Based on Swap-VAE [19], we modify the assumption of independence along time and introduce state variables to accumulate and filter temporal information, so that we can construct highly time-dependent latent spaces. Our model is able to progressively compress neural activity along a chronological order in a natural way.
@@
  • Through evaluation on mouse visual datasets, we demonstrate that our model efficiently decodes neural activity to related natural scenes or natural movies, showing the highest performance compared to alternative models.
  • Through visualization of latent representations, we provide evidence that our model captures explicit temporal structures of neural dynamics for different time scales.
@@ 2 Related work
VAE for population neural data analysis
As the advancement of deep learning, VAE-based approaches have been proposed to discover latent variables underlying population neural activity, better elucidating the mechanisms of neural representation. Latent factor analysis via dynamical systems (LFADS) used RNNs in a modified sequential VAE framework, accurately predicting observed behavioral variables and extracting precise firing rate estimates on single trials from several motor cortical datasets [22, 16]. pi-VAE, incorporating the relationship between the latent and task variables, constructed interpretable and identifiable latent variables [36]. To learn disentangled latent neural representations without labels, Swap-VAE combined a generative modeling framework with self-supervised learning based on an alignment loss, building latent factors linked to behavior [19]. CEBRA, a self-supervised learning algorithm, obtained consistent embeddings of high-dimensional recordings using auxiliary variables [25].
Visual neural data analysis via latent variables
With the growth of simultaneously recorded neural population data from the visual cortex, latent variable methods have been applied to extract underlying manifolds of neural activity. For population activity in macaque V1, some work used GPFA and found that anesthesia-induced network state fluctuations lead to correlated variability [11], while another proposed nonlinear generative fLDS, capturing neural variability with a few latent dimensions under drifting grating stimulus [13]. A rectified latent variable model built latent variables that show a spectrum of functional groups as neurons in the murine visual cortex under drifting gratings [21]. [25] achieved rapid, high-accuracy decoding of natural videos.
3 Time-Dependent SwapVAE
@@
Considering the neural activity of a population of neurons over a period of time, we define a sequence input as $\mathbf{x}=\left(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T}\right)\in{\mathbb{R}^{{T}\times{N}}}$, which represents spike counts of $N$ neurons within $T$ small time windows. The output $\mathbf{r}=\left(\mathbf{r}_{1},\ldots,\mathbf{r}_{T}\right)\in{\mathbb{R}^{{T}\times{N}}}$ is an estimate of firing rates of the input. The low-dimensional latent representation of time point $t$ is denoted as $\mathbf{z}_{t}\in{\mathbb{R}^{M}}$.
@@ 3.1 Model Architecture
Our goal is to compress temporal neural activity into high-quality latent spaces with modeling the temporal relationship of sequences. Motivated by these considerations, we explicitly model the time dependence between latent variables across time steps on the basis of Swap-VAE [19]. Specifically, we divide the latent space into content space and style space ($\mathbf{z}_{t}=[\mathbf{z}_{t}^{(c)},\mathbf{z}_{t}^{(s)}]$), just like Swap-VAE. However, instead of assuming that the latent space is independent along the time dimension, we introduce the state variable $\mathbf{h}_{t}$ to sift and accumulate temporal information [2, 12, 5], so that the latent and output variables of the current time step are conditioned on the input and state variables of the subsequent time steps. Consequently, we name our model as Time-Dependent SwapVAE. The operation of each module in our model is shown in Figure 1, and we formulate them below.
Figure 1: The method overview. A. The illustration of TiDe-SwapVAE for analyzing neural activity of the mouse visual cortex during passive viewing. The encoder of our model extracts features from sequential spike data. The latent variables are built conditioned on features from encoder and state variables of RNN to consider temporal dependence. The decoder maps latent variables to reconstructed firing rates. B. The illustrations of modules in our model. (see details in Section 3.1)
@@
As we add the state variables, the latent space is constructed not only from $\mathbf{x}_{t}$ but also from $\mathbf{h}_{t-1}$. While the content latent variables are constructed as deterministic values, the style latent variables are constructed as a parameterized conditional distribution (approximate posterior):
@@
where $f_{x}$, $f_{enc}^{(c)}$ and $f_{enc}^{(s)}$ are all parameter-learnable neural networks for extracting features from the input and computing the latent variables, respectively. The similar functions of $f$ in the following text are also trainable neural networks.
@@
Similar to the variational approximate posterior, the prior of $\mathbf{z}_{t}^{(s)}$ is conditioned on $\mathbf{h}_{t-1}$ for temporal dependence. The distribution is formulated as:
@@
The decoder aims to reconstruct the neural activity input, which receives the full latent variables and the state variables of the style space. Since the input is a sequence of spike counts, we denote the reconstructed responses as a parameterized Poisson distribution [22, 13], i.e., the actual output of the decoder is spike firing rates:
@@
The state variable is updated by recurrent neural networks, GRU [4]. Through selective integration and utilization of input and latent variables, the state variable is crucial for learning complex sequential dynamics. The updated functions are as follows:
@@ 3.2 Model Learning
We use self-supervised contrastive learning to jointly train the parameters of all modules in our model. For a given sample $\mathbf{x}=\left(\mathbf{x}_{1},\ldots,\mathbf{x}_{T}\right)$, we randomly select another sequence offset by several time steps as a positive sample, denoted $\mathbf{x}_{\mathrm{pos}}=\left(\mathbf{x}_{1+\Delta},\ldots,\mathbf{x}_{T+\Delta}\right)$, where $\Delta$ can be positive or negative. In this work, the offset is always less than the length of the sequence, ensuring that the positive sample pairs have overlaps. After building positive pairs, we follow the design of Swap-VAE. On the one hand, we are encouraged to bring the content latent variables of the positive pairs closer together. On the other hand, we swap the content latent variables between the positive pairs while maintaining the style space to construct swapped latent representations and use them to compute new reconstructed firing rates $\mathbf{\hat{r}}$ and $\mathbf{\hat{r}}_{\mathrm{pos}}$.
As we introduce time dependence into our model, we extend the evidence lower bound of VAE to a time-wise version and add the objective function for contrastive learning, which together form the loss function of TiDe-SwapVAE:
$\mathcal{L}=\mathcal{L}_{\mathrm{recons}}+\mathcal{L}_{\mathrm{swap}}+\beta\mathcal{L}_{\mathrm{regular}}+\gamma\mathcal{L}_{\mathrm{contrast}},$ (7)
where $\beta$ and $\gamma$ are hyperparameters that serve to control the severity of penalty for each loss terms. The reconstruction loss $\mathcal{L}_{\mathrm{recons}}$ and the swapped reconstruction loss $\mathcal{L}_{\mathrm{swap}}$ is time-wise, formulated as $\frac{1}{T}\sum_{t=1}^{T}{\left[\mathcal{L}_{\mathrm{P}}(\mathbf{x}_{t},\mathbf{r}_{t})+\mathcal{L}_{\mathrm{P}}(\mathbf{x}_{\mathrm{pos},t},\mathbf{r}_{\mathrm{pos},t})\right]}$ and $\frac{1}{T}\sum_{t=1}^{T}{\left[\mathcal{L}_{\mathrm{P}}(\mathbf{x}_{t},\mathbf{\hat{r}}_{t})+\mathcal{L}_{\mathrm{P}}(\mathbf{x}_{\mathrm{pos},t},\mathbf{\hat{r}}_{\mathrm{pos},t})\right]}$, respectively. $\mathcal{L}_{\mathrm{P}}$ is Poisson negative log likelihood loss. As for $\mathcal{L}_{\mathrm{contrast}}$, we do not apply time-wise operation, but flatten the temporal and spatial dimensions of the content latent variables from the positive pairs and compute the cosine distance between the two vectors. $\mathcal{L}_{\mathrm{regular}}$ is the KL divergence to measure the difference between the prior and the approximate posterior of the style latent variables, formulated as $\frac{1}{T}\sum_{t=1}^{T}{\left[\mathrm{D}_{\mathrm{KL}}(\mathbf{z}_{t}^{(s)}\|\mathbf{\tilde{z}}_{t}^{(s)})+\mathrm{D}_{\mathrm{KL}}(\mathbf{z}_{\mathrm{pos},t}^{(s)}\|\mathbf{\tilde{z}}_{\mathrm{pos},t}^{(s)})\right]}$. As described above, we assume that both distributions are Gaussian, so that the KL divergence can be written explicitly as:
@@
where all operations are element-wise calculations on vectors.
@@ 4.1 Evaluation and Alternative Models
In this work, we intend to compress neural activity to latent representations that exhibit strong relevance to the visual stimulation. According to this expectation, we evaluate our model from two aspects. First, we quantify the performance on decoding visual stimuli. Second, we assess the clarity of temporal structures extracted from neural dynamics.
For a comprehensive comparative analysis, we compare TiDe-SwapVAE with several latent variable models, including four generative models (an unsupervised: $\beta$-VAE [14], a sequential: LFADS [22], a supervised: pi-VAE [36], and a self-supervised: Swap-VAE [19]) and a nonlinear encoding method with contrastive learning (CEBRA) [25]. Specifically, $\beta$-VAE, pi-VAE and Swap-VAE compress neural activity independently for each time point. LFADS processes sequential neural activity with bidirectional RNNs. CEBRA encodes temporal features of sequence data with fixed convolutional kernels. None of them build the latent representations progressively along the chronological order.

### v2

# Time-Dependent VAE for Building Latent Representations from Visual Neural Activity with Complex Dynamics

@@ Abstract
Seeking high-quality representations with latent variable models (LVMs) to reveal the intrinsic correlation between neural activity and behavior or sensory stimuli has attracted much interest. Most work has focused on analyzing motor neural activity that controls clear behavioral traces and has modeled neural temporal relationships in a way that does not conform to natural reality. For studies of visual brain regions, naturalistic visual stimuli are high-dimensional and time-dependent, making neural activity exhibit intricate dynamics. To cope with such conditions, we propose Time-Dependent Split VAE (TiDeSPL-VAE), a sequential LVM that decomposes visual neural activity into two latent representations while considering time dependence. We specify content latent representations corresponding to the component of neural activity driven by the current visual stimulus, and style latent representations corresponding to the neural dynamics influenced by the organism’s internal state. To progressively generate the two latent representations over time, we introduce state factors to construct conditional distributions with time dependence and apply self-supervised contrastive learning to shape them. By this means, TiDeSPL-VAE can effectively analyze complex visual neural activity and model temporal relationships in a natural way. We compare our model with alternative approaches on synthetic data and neural data from the mouse visual cortex. The results show that our model not only yields the best decoding performance on naturalistic scenes/movies but also extracts explicit neural dynamics, demonstrating that it builds latent representations more relevant to visual stimuli.
@@ 1 Introduction
With the rapid development of neural recording technologies, researchers are now able to simultaneously record the spiking activity of large populations of neurons, providing new avenues for exploring the brain [39]. For analyzing these high-dimensional data, an important scientific problem is how to account for the intrinsic correlation between neural activity and behavioral patterns or sensory stimuli. As an influential approach, latent variable models (LVMs) construct low-dimensional latent representations bridging to behavior or stimuli and explain neural activity well [33, 2, 40, 21, 25]. Recently, advanced deep learning algorithms allowed LVMs to extract high-quality representations from neural activity without knowledge of experimental labels [42, 30, 17, 26], or to incorporate behavioral information into models to constrain the shaping of latent variables [27, 20, 32, 37]. These approaches have made various contributions to the analysis of neural activity, such as predicting held neural responses [16, 30], decoding related motion patterns or simple visual scenes [26, 34], and constructing interpretable latent structures [45, 1].
However, most studies have dealt with neural data recorded from motor brain areas under specific controlled behavioral settings [8, 30, 45, 26, 31], such as pre-planned reaching movements [13]. There is little work using LVMs to analyze neural data from visual brain regions [16, 44, 34], even though how the visual system encodes input to recognize objects is a primary topic [10], and decoding visual neural activity to visual stimuli is a challenging research highlight in the neuroscience community [22, 41, 12]. Furthermore, existing LVMs treat temporal relationships unnaturally [30, 34] or even don’t consider time dependence [45, 29]. Given that visual neural activity has strict antecedent time dependence, these models may struggle to build high-quality latent representations.
@@
In this work, we propose Time-Dependent Split VAE (TiDeSPL-VAE), a sequential LVM that builds two split latent representations with time dependence to better analyze visual neural activity. We adopt the practice of splitting latent variables into content and style representations [26]. Content latent representations correspond to the component of neural activity driven by the current visual stimulus, while style latent representations correspond to the neural dynamics influenced by the organism’s internal state (pupil position, signals relayed from other brain regions, the neurons’ underlying currents, etc.). These latent variables are optimized by self-supervised contrastive learning. For comparison with outstanding alternatives, we evaluate our model on synthetic and mouse visual datasets. The results show that our model builds meaningful latent representations that are highly correlated with complex visual stimuli, providing new insights into the intrinsic relationship between neural activity and visual stimulation. Specifically, our main contributions are as follows.
@@
  • To construct highly time-dependent latent representations, we introduce state factors to accumulate and filter temporal information, allowing TiDeSPL-VAE to progressively compress neural activity along a chronological order in a natural way. Besides, we apply self-supervised contrastive learning to enhance the time constraint and to shape latent variables.
  • Through evaluation on synthetic datasets, we show that our model better recovers latent structure and is good at handling time-sequential data.
@@
  • Through evaluation on mouse visual datasets, we demonstrate that our model decodes neural activity to related natural scenes or natural movies well, showing the highest performance compared to alternative models. Furthermore, visualization of latent representations presents that our model captures explicit temporal structures of neural dynamics for different time scales.
@@ 2 Related Work
With the advancement of deep learning, the application of cutting-edge learning algorithms and the innovative design of model structures have greatly promoted the development of LVMs in neuroscience. Some prominent works are summarized below.
VAE-based LVMs for neural activity analysis
Recently, VAE-based approaches have become a major avenue to discover latent variables underlying population neural activity, which better elucidates the mechanisms of neural representations. As a well-known model, latent factor analysis via dynamical systems (LFADS) used RNNs in a sequential VAE framework, extracting precise firing rate estimates and predicting observed behavior for single-trial data on motor cortical datasets [30, 23, 24]. Through specific latent variable design, pi-VAE [45] and Swap-VAE [26] built interpretable latent structures linked to motor behavioral patterns.
LVMs for visual neural activity analysis
Several studies have made an effort to extract latent manifolds from visual neural activity using LVMs. Although these studies cover various types of models, such as the Gaussian process model [14, 44, 18], linear dynamical system [16], and autoencoder [29], they are limited to simple visual stimuli such as gratings and are used only for the downstream task of reconstructing neural responses. Recently, CEBRA, a self-supervised learning model, obtained consistent latent representations and made progress in decoding movies [34].
3 Time-Dependent Split VAE
@@
Considering the neural activity of a population of neurons over a period of time, we define a sequence input as $\mathbf{x}=\left(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T}\right)\in{\mathbb{R}^{{T}\times{N}}}$, which represents spike counts of $N$ neurons within $T$ time windows. The output $\mathbf{r}=\left(\mathbf{r}_{1},\ldots,\mathbf{r}_{T}\right)\in{\mathbb{R}^{{T}\times{N}}}$ is an estimate of firing rates of the input. The low-dimensional latent representation of time point $t$ is denoted as $\mathbf{z}_{t}\in{\mathbb{R}^{M}}$.
@@ 3.1 Model Architecture
Our goal is to compress neural activity into high-quality latent representations while modeling the temporal relationship. Therefore, we explicitly model the time dependence between latent variables across time steps based on a sequential VAE. In practice, we split latent variables into content and style latent representations ($\mathbf{z}_{t}=[\mathbf{z}_{t}^{(c)},\mathbf{z}_{t}^{(s)}]$), corresponding to stimulus-driven and neural dynamical components, respectively. To construct time-dependent connections along a chronological order, we introduce the state factor $\mathbf{h}_{t}$ to sift and accumulate temporal information [4, 15, 7] so that the latent variables and output of the current time step are conditioned on the input and state factors of the antecedent time steps. We name our model as Time-Dependent Split VAE (TiDeSPL-VAE; Figure 1A) and formulate the operations of each module (Figure 1B) below.
Figure 1: The method overview. A. The illustration of TiDeSPL-VAE for analyzing neural activity in the mouse visual cortex during passive viewing. The encoder extracts spatial features from sequential spike data. The latent variables are built conditionally on features of the encoder and RNNs’ state factors to introduce time dependence. The decoder maps latent variables to reconstructed firing rates. Detailed network structures are given in Appendix A. B. The operations of each module in TiDeSPL-VAE (see details in Section 3.1).
@@
Based on the above assumptions, content latent variables driven by the current stimulus are constructed as deterministic values, while style latent variables are constructed as random values from a parameterized distribution (approximate posterior) since there is a lot of intrinsic noise as well as variability in the neural dynamics. For time dependence, the latent variables are built on $\mathbf{x}_{t}$ and $\mathbf{h}_{t-1}$:
@@
where $f_{x}$, $f_{enc}^{(c)}$ and $f_{enc}^{(s)}$ are all parameter-learnable neural networks for extracting spatio-temporal features and building the latent variables. The similar functions of $f$ in the following text are also trainable neural networks.
@@
Similar to the variational approximate posterior, the prior of $\mathbf{z}_{t}^{(s)}$ is conditioned on $\mathbf{h}_{t-1}$ for time dependence. The distribution is formulated as:
@@
The decoder aims to reconstruct the neural activity input, which receives the full latent variables with style state factors as an auxiliary. Since the input is a sequence of spike counts, we denote the reconstructed responses as a parameterized Poisson distribution [16], i.e., the actual output of the decoder is spike firing rates:
@@
The state factor is updated by recurrent neural networks, GRU [6]. By selectively integrating and exploiting input and latent variables, the state factor is crucial for learning complex sequential dynamics. Importantly, since content state factors depend on dynamic changes in visual stimuli while stimulus-driven neural activity inevitably affects neural dynamics, $\mathbf{h}_{t}^{(c)}$ and $\mathbf{h}_{t}^{(s)}$ are updated differently:
@@ 3.2 Model Learning
We use self-supervised contrastive learning to jointly train the parameters of all modules in our model. For a given sample $\mathbf{x}=\left(\mathbf{x}_{1},\ldots,\mathbf{x}_{T}\right)$, we randomly select another sequence offset by several time steps as a positive sample, denoted $\mathbf{x}_{\mathrm{pos}}=\left(\mathbf{x}_{1+\Delta},\ldots,\mathbf{x}_{T+\Delta}\right)$, where $\Delta$ can be positive or negative. In this work, the offset is always less than the length of the sequence to ensure that the positive sample pairs overlap. Then, a mini-batch of negative samples is randomly selected from the entire training set. The model is encouraged to bring the content latent variables of the positive pairs closer together and to push those of negative samples away.
As we introduce time dependence into our model, we extend the evidence lower bound of VAE to a time-wise version and add the objective function for contrastive learning, which together form the loss function of TiDeSPL-VAE:
$\mathcal{L}=\mathcal{L}_{\mathrm{recons}}+\beta\mathcal{L}_{\mathrm{regular}}+\gamma\mathcal{L}_{\mathrm{contrast}},$ (7)
where $\beta$ and $\gamma$ are hyperparameters that serve to control the severity of the penalty for each loss term. The reconstruction loss $\mathcal{L}_{\mathrm{recons}}$ is formulated as $\frac{1}{T}\sum_{t=1}^{T}{\left[\mathcal{L}_{\mathrm{P}}(\mathbf{x}_{t},\mathbf{r}_{t})+\mathcal{L}_{\mathrm{P}}(\mathbf{x}_{\mathrm{pos},t},\mathbf{r}_{\mathrm{pos},t})\right]}$, where $\mathcal{L}_{\mathrm{P}}$ is Poisson negative log likelihood loss. $\mathcal{L}_{\mathrm{regular}}$ is the KL divergence to measure the difference between the prior and the approximate posterior of style latent variables, formulated as $\frac{1}{T}\sum_{t=1}^{T}{\left[\mathrm{D}_{\mathrm{KL}}(\mathbf{z}_{t}^{(s)}\|\mathbf{\tilde{z}}_{t}^{(s)})+\mathrm{D}_{\mathrm{KL}}(\mathbf{z}_{\mathrm{pos},t}^{(s)}\|\mathbf{\tilde{z}}_{\mathrm{pos},t}^{(s)})\right]}$. As described above, we assume that both distributions are Gaussian, so that the KL divergence can be written explicitly as:
@@
where all operations are element-wise calculations on vectors. Besides, we added L2 regularization to the expectation and log-variance of the prior distribution to stabilize model training. We utilize NT-Xent loss [5] as $\mathcal{L}_{\mathrm{contrast}}$. For this term, we do not apply the time-wise operation, but flatten the temporal and spatial dimensions of content latent variables for the loss computation. To enhance the effect of the positive sample, we adopt the practice of swapping content latent variables between the positive pairs while maintaining style latent variables [26]. The swapped latent representations are then used to compute new reconstructed firing rates and an additional reconstruction loss. A detailed derivation of all formulas is given in Appendix B.
@@ 4.1 Evaluation and Alternative Models
In this work, we intend to construct latent representations that exhibit strong relevance to visual stimulation. According to this expectation, we evaluate our model from two aspects, similar to studies oriented to motor brain regions [26, 34]. First, we quantify the performance in decoding visual stimuli using latent representations, which has long served as a research hotspot for unraveling the mechanisms of visual processing [22, 41]. Second, we assess the clarity of latent temporal structures extracted from neural dynamics.
For a comprehensive analysis, we compare TiDeSPL-VAE with several outstanding LVMs, including four generative models (an unsupervised: $\beta$-VAE [19], a sequential: LFADS [30], a supervised: pi-VAE [45], and a self-supervised: Swap-VAE [26]) and a nonlinear encoding method with contrastive learning (CEBRA) [34]. Specifically, $\beta$-VAE, pi-VAE and Swap-VAE compress neural activity independently for each time point. LFADS processes sequential neural activity with bidirectional RNNs. CEBRA encodes temporal features of sequence data with fixed convolutional kernels. None of them build latent representations progressively along the chronological order.